OpenStudy (anonymous):

Help solve 4x^3-12x^2-x+3=0? 1) list all possible rational roots? 2) solve the equation. Solve the equation x^3-8x^2+21x-20=0 given that one of the roots is 2+i

OpenStudy (tkhunny):

Well, do it. Factors of 3 divided by factors of 4. That Is ALL the possible Rational Roots. With Real coefficients, Complex Roots must come in pairs. Consider 2-i, also.

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

Ok, you can get 4x^3-12x^2-x+3=0 into this form 4x^2(x-3)-(x-3)=0 4x^2(x-3)=x-3 4x^2=1 x^2=1/4 x=+- 1/2 real roots are +1/2 and -1/2

OpenStudy (anonymous):

@cblrtopas is the 3rd root 3?

OpenStudy (dan815):

u know how to do synthetic division?

OpenStudy (anonymous):

Yes actually, I think it is

OpenStudy (anonymous):

it satisfies the equation so yes....

OpenStudy (anonymous):

@dan815 yes @cblrtopas okay, but i dont know how 3 was found. i didnt do that, someone told me.

OpenStudy (anonymous):

@cblrtopas and also, when you worked out the steps, you put 4x^2 and its to the 3rd power.

OpenStudy (dan815):

well you look at the coefficient of 1st and last numbers so 4x^3-12x^2-x+3=0? so 4 and 3 you try multiples so 1 , 2 , 3 , 4 you find 3 works

OpenStudy (anonymous):

I eliminated x=3 from the solution set by dividing by x-3 in my derivation earlier. When dividing by x-3 I was implicitly assuming that the quantity x-3 was not zero. In this case it's possible that x was indeed 3. Basically, you can get the third root by using a cubic equation or some fancy factorization. I don't remember this crap from the top of my head I need to look at my algebra book.

OpenStudy (dan815):

then (x-3) syenthtic division gives you (x-3)(2x-1)(2x+1)

OpenStudy (anonymous):

tessa11, you need to go through each step carefully. The 4x^2 is multiplied by (x-3). which will yield the cubic when multiplied out. Don't copy my work make the derivation yourself..

OpenStudy (anonymous):

@cblrtopas thank you! i think i get it, but how would i do the 2nd equation? because it gives me a root, and it has an imaginary number in it.

OpenStudy (anonymous):

what second equation?

OpenStudy (anonymous):

@cblrtopas Solve the equation x^3-8x^2+21x-20=0 given that one of the roots is 2+i

OpenStudy (anonymous):

You didn't copy the problem correctly I think (2-i) is not a root if there's no x somewhere in it.

OpenStudy (anonymous):

@cblrtopas no thats what my teacher gave me exactly :(

OpenStudy (anonymous):

Ugh......It's late.....do you even know what imaginary numbers are used for... Seriously screw imaginary algebra .....it will never become useful I don't know girl...

OpenStudy (anonymous):

@cblrtopas haha alright, well thank you for your help!!! :)

OpenStudy (anonymous):

\[(x-3) (2 x-1) (2 x+1)=0 \]